# The stochastic gravitational wave background from cosmic superstrings

### v1

@@ IV.4 Forecast for future experiments
Gravitational waves in the nanohertz range are currently being probed by NANOGrav Agazie et al. (2023), with low-frequency ranges beyond current PTAs being within the sensitivity ranges of SKA Janssen et al. (2015) and ROMAN Wang et al. (2022). The microhertz to hertz range will be covered by LISA Bartolo et al. (2016); Caprini et al. (2019); Auclair et al. (2023) and cold-atom-based experiments AEDGE El-Neaj et al. (2020) and AION Badurina et al. (2020). This range could be useful in differentiating between kink and cusp sourced GW emission, as the fits to NG15 data differ significantly in magnitude within this range. Finally, the highest frequency range ($\sim 10-100$ Hz) of the spectrum could already be within the sensitivity range of LIGO/Virgo/KAGRA Aasi et al. (2015); Abbott et al. (2016, 2019)(especially when they reach their maximal sensitivities), although drop-offs in magnitude could be caused by modifications to standard cosmic evolution Ellis et al. (2023). Due to its larger magnitude at high frequencies, the large-loop spectrum will be easiest to probe by current and proposed future experiments. As the small-loop spectrum has a sharper drop from peak emission intensity, its high-frequency tail is beyond the reach of LIGO/Virgo/KAGRA (but within the projected sensitivity of the Einstein Telescope Hild et al. (2011); Punturo et al. (2010)).
@@ V Conclusions
Fitting the cosmic superstring model to the NANOGrav 15-year gravitational wave background data gives well-constrained values for the fundamental string tension at both small and large loop sizes. As string loops oscillate, the dominant fraction of gravitational waves can come from either cusps (single points moving instantaneously at highly relativistic speeds) or kinks (discontinuities in the loop tangent vector). Within the NANOGrav frequency range, this origin does not have a significant effect on the best-fit spectra or parameters, as we obtain expectation values for the fundamental string tension of $\log_{10}(G\mu_{1})=-11.5^{+0.3}_{-0.3}$($-11.6^{+0.2}_{-0.3}$) for gravitational waves originating from large cuspy (kinky) cosmic superstring loops and $\log_{10}(G\mu_{1})=-9.7^{+0.7}_{-0.7}$($-9.9^{+1.0}_{-0.5}$) for small cuspy (kinky) loops. The posteriors for string coupling, $g_{s}$, and $w$, which describes the size of extra compact dimensions, are much flatter and we set $2\sigma$ confidence bounds as $g_{s}<0.65$($<0.65$), $w<0.91$($<0.92$) for the large cuspy (kinky) loops and $g_{s}<0.63$($<0.61$), $w<0.88$($<0.83$) for small cuspy (kinky) loops. Finally, we repeat the analysis by assuming the string loops consist of 10% large loops and 90% small loops, where only large loops contribute to the GW signal. In this case, we obtain bounds for the string tension consistent with Afzal et al. (2023) and Figueroa et al. (2024), who make the same assumptions. We also place $2\sigma$ bounds $g_{s}<0.54$ and $w<0.82$, which is not possible with the single-string approximation used in Afzal et al. (2023); Figueroa et al. (2024). In all cases, the quality of fits is comparable to the current best astrophysical models involving supermassive black hole binaries, although large loops are favoured over small loops.
@@
There remain significant uncertainties in modelling the GW signal from cosmic (super)string networks. The most important unknown is the size of string loops, which determines both the magnitude and shape of the spectrum as seen, for example, in Figs. 2 and 3. This is a challenging and important open problem in the field. The best we can do at present is model various options, through different choices of loop distributions, or the parameter $\alpha$ as we have done in this paper. For cosmic superstrings, in particular, there are additional uncertainties arising from approximations/assumptions in the computation of the self- and cross-string interaction parameters $\tilde{c}_{i}$ and $\tilde{d}^{k}_{ij}$ in the multi-tension VOS model, the impact of compactification choices for the extra dimensions, and other assumptions relating to the formation and evolution of string junctions (see for example Rybak et al. (2019)). Nevertheless, our current modelling techniques provide a systematic way of quantifying these effects as we have exemplified in this paper. Indeed, given these modelling uncertainties, the relative robustness of the predicted GW signal is somewhat surprising. In particular, a single species toy-model can provide a reasonable approximation to the GW spectrum over most of the relevant frequency range. The detailed modelling we have discussed here, crucially, links the GW spectrum with the underlying fundamental parameters of the model – the tension of the fundamental string $\mu_{F}$, the string coupling constant $g_{s}$ and the size of the extra dimensions through $w$ – and makes the characteristic features of cosmic superstring networks apparent at low frequencies. This will be important for constraining or detecting these effects with future GW data.
@@ Acknowledgements.
The work of AA, EJC and AM was supported by an STFC Consolidated Grant [Grant No. ST/X000672/1]. JR was supported by an STFC studentship [Grant No. ST/Y509437/1]. We are grateful to Lara Sousa for helpful discussions. For the purpose of open access, the authors have applied a CC BY public copyright license to any Author Accepted Manuscript version arising.

### v2

@@ IV.4 Forecast for future experiments
Gravitational waves in the nanohertz range are currently being probed by NANOGrav Agazie et al. (2023), with low-frequency ranges beyond current PTAs being within the sensitivity ranges of SKA Janssen et al. (2015) and ROMAN Wang et al. (2022). The microhertz to hertz range will be covered by LISA Bartolo et al. (2016); Caprini et al. (2019); Auclair et al. (2023) and cold-atom-based experiments AEDGE El-Neaj et al. (2020) and AION Badurina et al. (2020). This range could be useful in differentiating between kink and cusp sourced GW emission, as the fits to NG15 data differ significantly in magnitude within this range. Finally, the highest frequency range ($\sim 10-100$ Hz) of the spectrum could already be within the sensitivity range of LIGO/Virgo/KAGRA Aasi et al. (2015); Abbott et al. (2016, 2019)(especially when they reach their maximal sensitivities), although drop-offs in magnitude could be caused by modifications to standard cosmic evolution Ellis et al. (2023); Datta and Samanta (2025). Similar reductions in high-frequency magnitude could also be caused by explicit modelling of gravitational back-reaction on cosmic superstring loops Wachter et al. (2024a). Due to its larger magnitude at high frequencies, the large-loop spectrum will be easiest to probe by current and proposed future experiments. As the small-loop spectrum has a sharper drop from peak emission intensity, its high-frequency tail is beyond the reach of LIGO/Virgo/KAGRA (but within the projected sensitivity of the Einstein Telescope Hild et al. (2011); Punturo et al. (2010)).
@@ V Conclusions
Fitting the cosmic superstring model to the NANOGrav 15-year gravitational wave background data gives well-constrained values for the fundamental string tension at both small and large loop sizes. As string loops oscillate, the dominant fraction of gravitational waves can come from either cusps (single points moving instantaneously at highly relativistic speeds) or kinks (discontinuities in the loop tangent vector). Within the NANOGrav frequency range, this origin does not have a significant effect on the best-fit spectra or parameters, as we obtain expectation values for the fundamental string tension of $\log_{10}(G\mu_{1})=-11.4^{+0.3}_{-0.2}$($-11.5^{+0.3}_{-0.2}$) for gravitational waves originating from large cuspy (kinky) cosmic superstring loops and $\log_{10}(G\mu_{1})=-9.7^{+0.7}_{-0.7}$($-9.9^{+1.0}_{-0.5}$) for small cuspy (kinky) loops. The posteriors for string coupling, $g_{s}$, and $w$, which describes the size of extra compact dimensions, are much flatter and we set $2\sigma$ confidence bounds as $g_{s}<0.69$($<0.70$), $w<0.90$($<0.91$) for the large cuspy (kinky) loops and $g_{s}<0.63$($<0.61$), $w<0.88$($<0.83$) for small cuspy (kinky) loops.
Finally, we repeat the analysis by assuming the string loops consist of 10% large loops and 90% small loops, where only large loops contribute to the GW signal. In this case, we obtain bounds for the string tension of $log_{10}(G\mu_{1})^{NG}=-11.0^{+0.4}_{-0.2}$($-11.1^{+0.4}_{-0.3}$) which are different from the bounds obtained by Afzal et al. (2023) and Figueroa et al. (2024), who make the same assumptions on the string loop distribution. These differences can be attributed to our modelling of the cosmic superstring network in terms of multiple correlation lengths, as opposed to the single-string approximation used in Afzal et al. (2023); Figueroa et al. (2024). As discussed in section IV.3, calculating the superstring GW spectrum by rescaling an ordinary cosmic string spectrum can lead to loop lengths larger than the string correlation length. We also place $2\sigma$ bounds on $g_{s}<0.59$$(<0.61)$ and $w<0.84$($<0.84$), which is not possible with the single-string approximation used in Afzal et al. (2023); Figueroa et al. (2024). In all cases, the quality of fits is comparable to the current best astrophysical models involving supermassive black hole binaries, although large loops are favoured over small loops.
@@
There remain significant uncertainties in modelling the GW signal from cosmic (super)string networks. The most important unknown is the size of string loops, which determines both the magnitude and shape of the spectrum as seen, for example, in Figs. 2 and 3. This is a challenging and important open problem in the field. The best we can do at present is model various options, through different choices of loop distributions, or the parameter $\alpha$ as we have done in this paper. For cosmic superstrings, in particular, there are additional uncertainties arising from approximations/assumptions in the computation of the self- and cross-string interaction parameters $\tilde{c}_{i}$ and $\tilde{d}^{k}_{ij}$ in the multi-tension VOS model, the impact of compactification choices for the extra dimensions, and other assumptions relating to the formation and evolution of string junctions (see for example Rybak et al. (2019)). Nevertheless, our current modelling techniques provide a systematic way of quantifying these effects as we have exemplified in this paper. Indeed, given these modelling uncertainties, the relative robustness of the predicted GW signal is somewhat surprising. In particular, a single species toy-model can provide a reasonable approximation to the GW spectrum over most of the relevant frequency range, as heavier strings only contribute significantly at low frequencies. If the single-string approximation is made, however, it is important to directly model the dependence of correlation length on intercommuting probability in order to avoid unphysical loop sizes.
The detailed modelling we have discussed here, crucially, links the GW spectrum with the underlying fundamental parameters of the model – the tension of the fundamental string $\mu_{F}$, the string coupling constant $g_{s}$ and the size of the extra dimensions through $w$ – and makes the characteristic features of cosmic superstring networks apparent at low frequencies. This will be important for constraining or detecting these effects with future GW data.
@@ Acknowledgements.
The work of AA, EJC and AM was supported by an STFC Consolidated Grant [Grant No. ST/X000672/1]. JR was supported by an STFC studentship [Grant No. ST/Y509437/1]. We are grateful to Lara Sousa and Ivan Rybak for helpful discussions. For the purpose of open access, the authors have applied a CC BY public copyright license to any Author Accepted Manuscript version arising.